# A Dimca-Greuel type inequality for foliations

### v3

@@ Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$. We prove the following Dimca-Greuel type inequality $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$, $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$, and $GSV_{p}(\mathcal{F},B)$ is the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ with respect to $B$. As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes, but it is in line with the idea developed by Wang.
@@
Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, Dimca-Greuel bound
@@ 1. Introduction
Let $\mathbb{C}\{x,y\}$ be the ring of complex convergent power series in two variables. Consider the germ of isolated plane curve singularity $C$ defined by the non-unit $f(x,y)\in\mathbb{C}\{x,y\}$. The Milnor number of $C$ is
$\mu(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(\partial_{x}f,\partial_{y}f)$
and its Tjurina number is
$\tau(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(f,\partial_{x}f,\partial_{y}f).$
Is clear from their definitions that $\tau(C)\leq\mu(C)$.
In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity. The question was affirmatively answered in [1], [10], and [17] for the irreducible case and in [2] for the reduced case.
It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs. More precisely, if $\mathcal{F}$ is a germ of singular holomorphic foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}\{x,y\}/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$. Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9, p. 4], [11, Definition 3.1]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension
$\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,Q,f).$
If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
(1) $\frac{\mu_{p}(\mathcal{F})}{\tau_{p}(\mathcal{F},\mathcal{B}_{0})}$
equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$. If the foliation $\mathcal{F}$ is of the second type (see [9, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied). It is natural to wonder if we can improve the bound $2$ for the ratio (1).
We note that, in general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
$\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
On the other hand, following Camacho-Lins Neto-Sad [4, Section 4], we can consider the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition). We note that if $B=\{f(x,y)=0\}$, and $\mathcal{F}:df=0$, then $\mu_{p}(\mathcal{F},B)$ also coincides with the Milnor number $\mu(B)$ of $B$ (see Remark 2.1). Therefore, we could think to studying if the inequality
$03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0
holds. But, for the germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},0)$ defined by
(2) $\omega=y\left(13x^{7}+6y^{7}\right)dx-x\left(5y^{7}+12x^{7}\right)dy,$
and the separatrix
(3) $B:x^{13}+x^{6}y^{7}+y^{12}=0,$
we have $\mu_{0}(\mathcal{F},B)=85$ and $\tau_{0}(\mathcal{F},B)=61$. Thus
$113\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61=255-244=$ 11 .
In order to establish an inequality of the Dimca-Greuel’s type for foliations, it is necessary to add the Gómez-Mont–Seade–Verjovsky index (in short GSV-index) (see [12]) as stating our main result.
@@
(4) $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$
@@
We will prove Theorem A in Section 3. For the foliation given in (2), and $B$ as in (3), we have $GSV_{0}(\mathcal{F},B)=-47$. Hence,
$03\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)+GSV_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61+(-47)=-36<$ 0
verifying Theorem A.
As a consequence of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17] (see Corollary 5.1).
@@ 2. Multiplicity of a foliation along a divisor of separatrices
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$).
@@
Consequently when $B$ is quasihomogeneous by ([15], Satz p.123 a) and d)] and (8) or by [18, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
@@ 3. Proof of Theorem A
Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$). Put $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$. If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([17]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$. By [17, Claim 4.4] we have
@@
From [17, Equation (4)] we obtain
@@ 3.1. Proof of Theorem A
where $B^{(i+1)}$ denotes the strict transform of $B^{(i)}$ thought $q^{(i)}$ for each $i=0,\ldots,N-1$, $B^{(N)}$ is smooth, and $\mathcal{F}^{(N)}$ is a foliation with a simple singularity at $q^{(N)}\in B^{(N)}$, or $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$, and $B^{(N)}$ is a dicritical separatrix (cf. [11, p. 1423]). Put $\mu^{(i)}:=\mu_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ for any $i\in\{1,\ldots,N\}$.
@@ 5. Applications
From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17]:
Corollary 5.1.
Let $f(x,y)\in\mathbb{C}\{x,y\}$ be an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
Proof.
It is a consequence of Remark 2.1 and applying Theorem A to the foliation $\mathcal{F}:df=0$, because $\mu_{p}(\mathcal{F},C)=\mu(C)$, $\tau_{p}(\mathcal{F},C)=\tau(C)$, and $GSV_{p}(\mathcal{F},C)=0$. ∎
Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator of (1) the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
@@
(21) $03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0
@@
$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})=3\cdot 25-4\cdot 26=-29<$ 0 .

### v4

# An upper bound for the GSV-index of a foliation

@@ Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and let $B$ be a separatrix of $\mathcal{F}$. We prove the following upper bound $GSV_{p}(\mathcal{F},B)\leq 4\tau_{p}(\mathcal{F},B)-3\mu_{p}(\mathcal{F},B)$, where $GSV_{p}(\mathcal{F},B)$ is the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ with respect to $B$, $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$.
@@
Holomorphic foliations, Gómez-Mont-Seade-Verjovsky index, Tjurina number, Multiplicity of a foliation along a divisor of separatrices.
@@ 1. Introduction
Let $\mathcal{F}$ be a germ of singular holomorphic foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, where $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$, and $B$ be a separatrix of $\mathcal{F}$. Several numerical invariants can be studied for the pair $(\mathcal{F},B)$, such as the Gómez-Mont–Seade–Verjovsky index (see [10]) of $\mathcal{F}$ with respect to $B$, denoted by $GSV_{p}(\mathcal{F},B)$, the multiplicity of $\mathcal{F}$ along $B$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and the Tjurina number of $\mathcal{F}$ with respect to $B:f(x,y)=0$, defined by
$\tau_{p}(\mathcal{F},B):=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,Q,f).$
In this paper, inspired by the blow-up techniques developed by Wang [14], we establish an optimal upper bound for $GSV_{P}(\mathcal{F},B)$ in terms of $\mu_{p}(\mathcal{F},B)$ and $\tau_{p}(\mathcal{F},B)$ as follows:
@@
(1) $GSV_{p}(\mathcal{F},B)\leq 4\tau_{p}(\mathcal{F},B)-3\mu_{p}(\mathcal{F},B)$
@@
Note that if $B=\{f=0\}$ is an irreducible plane curve germ, then by applying Theorem A to the foliation $\mathcal{F}:df=0$, we obtain the bound stated in [6, Question 4.2] and proved simultaneously in [1], [8] and [14], since $\mu_{p}(\mathcal{F},B)=\mu(B)$, $\tau_{p}(\mathcal{F},B)=\tau(B)$, and $GSV_{p}(\mathcal{F},B)=0$ in this case. Here, $\mu(B)$ and $\tau(B)$ denote the classical Milnor and Tjurina numbers of $B$. A complete answer to this question was given by Almirón [2]. Moreover, the author proposed a broader perspective of the study of the difference between Milnor and Tjurina numbers for isolated complete intersection singularities.
@@ 2. Multiplicity of a foliation along a divisor of separatrices
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [11, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$).
@@
Consequently when $B$ is quasihomogeneous by ([12], Satz p.123 a) and d)] and (5) or by [15, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
@@ 3. Proof of Theorem A
Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$). Put $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$. If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([14]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$. By [14, Claim 4.4] we have
@@
From [14, Equation (4)] we obtain
@@ 3.1. Proof of Theorem A
where $B^{(i+1)}$ denotes the strict transform of $B^{(i)}$ thought $q^{(i)}$ for each $i=0,\ldots,N-1$, $B^{(N)}$ is smooth, and $\mathcal{F}^{(N)}$ is a foliation with a simple singularity at $q^{(N)}\in B^{(N)}$, or $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$, and $B^{(N)}$ is a dicritical separatrix (cf. [9, p. 1423]). Put $\mu^{(i)}:=\mu_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ for any $i\in\{1,\ldots,N\}$.
@@ 5. Applications
We define the sum $T_{p}(\mathcal{F},C):=\sum_{B\subset C}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality for foliations:
@@
(18) $GSV_{p}(\mathcal{F},\mathcal{B})<4T_{p}(\mathcal{F},\mathcal{B})-3\mu_{p}(\mathcal{F},\mathcal{B})$
@@
$GSV_{p}(\mathcal{F},\mathcal{B})=0<4T_{p}(\mathcal{F},\mathcal{B})-3\mu_{p}(\mathcal{F},\mathcal{B})=29.$